OpenStudy (anonymous):

A train can travel 50% faster than a car. Both start from point A at the same time and reach point B 75 kms away from A at the same time. On the way, however, the train lost about 12.5 minutes while stopping at the stations. What is the speed of the car?

OpenStudy (anonymous):

trains speed=dist/time so 50/100x=75/y-12.5 x=speed of car and y=time

OpenStudy (anonymous):

car speed is x=75/y

OpenStudy (anonymous):

sorry 12.5minutes so/60

Miracrown (miracrown):

x = speed of car (km / min) What would the speed of the train be in km/min ?

OpenStudy (anonymous):

x/2=75/y-12.5/60

Miracrown (miracrown):

Not quite. The train goes 50% faster than the car the car goes "x" km/min - the train is going faster so, we can represent the speed of the train as 1.5x

Miracrown (miracrown):

1.5x = speed of train (km / min)

Miracrown (miracrown):

Now, have you tried to draw a diagram of this problem situation ? O.o

Miracrown (miracrown):